Lindiwe decides that she would like to buy her daughter, Mbali, a car when she turns 21 in six years' time. She deposits R6 000 each month into an account earning 8,94% interest per year, compounded monthly. The amount that Lindiwe (rounded to the nearest rand) will have available six years from now is


1
Expert's answer
2022-03-07T08:34:02-0500

According to compounded interest formula

"R=6000*(1+{\frac {0.0894} {12}})^{12*6}=6000*1.00745^{72}\approx 10239" Lindwie will have available six years from now